OpenStudy (anonymous):

-2d-1.7=3.9 how would I work out this problem

OpenStudy (anonymous):

@AlexandervonHumboldt2

rvc (rvc):

u already did a similar problem try it out

OpenStudy (welshfella):

first add 1.7 to both sides -2d - 1.7 + 1.7 = 3.9 + 1.7 -2d = 5.6 now divide both sides by -2 and you'll get the values of d